How many gallons of 65% antifreeze solution and a 35% antifreeze solution must be mixed to obtain 15 gallons of 53% antifreeze solution?
Answer by mananth(16946)

percent percentage ---------------- quantity
antifreeze I 65 ---------------- x gal
antifreeze II 35 ---------------- 15 - x gal
Mixture 53.00% ---------------- 15

65 x + 35 ( 15 - x ) = 53 * 15

65 x + 525 - 35 x = 795
65 x - 35 x = 795 - 525
30 x = 270
/ 30
x= 9 gal 65.00% antifreeze I
6 gal 35.00% antifreeze II